OpenStudy (anonymous):

Multiply (3.3 × 104) × (3 × 1018). Express the answer in scientific notation

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what.?

OpenStudy (anonymous):

@kaliyahfrank

OpenStudy (anonymous):

9.9x1022

OpenStudy (anonymous):

(3.3 * 10^4) * (3 * 10^18) = (3.3 * 3) * (10^4 * 10^18) = 9.9 * 10^(4 + 18) = 9.9 * 10^22

OpenStudy (tester97):

Thank you for showing your work instead of just giving the answer @cutepochacco Most people just give out the answer. Keep up the good work!